# Trajectory Prediction for Autonomous Driving: Progress, Limitations, and Future Directions

### v2

@@ 4.2 Injecting Driving Knowledge
Formalizing and incorporating driving knowledge into prediction models can enhance AVs reasoning and adaptability to uncertain or unfamiliar scenarios. Human drivers accumulate this knowledge through learning and experience. Similarly, AVs can benefit from a combination of injected driving knowledge and learned patterns from data. Table 14 provides a brief summary of learning-based models, which incorporate driving knowledge elaborating on how the knowledge is injected and how it is formalized.
@@
For agent-related knowledge, several works focus on integrating kinematic parameters to ensure physically feasible trajectory predictions. Xie et al. [218] proposed the Interactive Multiple Model (IMM) method, which leverages vehicle kinematic parameters such as acceleration and velocity in a physics-based model. Cui et al. [488] embedded a kinematic model of two-axle vehicles into the prediction pipeline, constraining outputs like turning radius, acceleration, and heading to align with vehicle dynamics. Additionally, works [490, 481] incorporated vehicle dynamics and kinematics constraints through Model Predictive Control (MPC), ensuring that predicted trajectories comply with speed limits, acceleration, and other physical constraints.
@@
In terms of traffic-related knowledge, DESIRE [64] incorporates this knowledge through a predefined cost function within its Inverse Optimal Control (IOC)-based ranking module. The cost function evaluates trajectory predictions based on criteria such as collision avoidance, road compliance, and interaction dynamics, assigning lower costs to socially and contextually plausible trajectories to ensure alignment with expected traffic norms. Greer et al. [486] proposed the YawLoss framework, which integrates traffic-related knowledge by enforcing vehicle motion constraints aligned with lane directionality. This is achieved through a designed auxiliary loss function, YawLoss, which penalizes trajectories deviating from the expected lane heading, thereby ensuring adherence to legal and social driving norms, such as lane-following behavior. Cho et al. [309] integrates traffic-related knowledge by representing predefined traffic rules as Signal Temporal Logic (STL) formulas. The model jointly predicts future trajectories and the degree of satisfaction (robustness slackness) for these rules, filtering out invalid trajectory predictions for surrounding vehicles.
@@
Map-related knowledge represents the largest class of models integrating driving knowledge. Wang et al. [491] proposed a lane-based prediction model that represents lane segments as vectorized elements, encoding geometric and semantic information such as centerlines, lane types, and heading angles, enabling interaction modelling between agents and lane segments. Deo et al. [480] integrates map-related knowledge by representing HD maps as lane graphs, capturing geometric and semantic details like lane centerlines, stop lines, crosswalks, and traffic flow directions. Nodes represent lane segments, while edges encode successor relationships and potential lane changes, ensuring trajectories align with legal routes. Yoon et al. [487] proposed a road-aware prediction model that incorporates curvilinear coordinates, lane boundaries, and feasible maneuvers from HD maps, using this information as prior knowledge to constrain trajectory predictions. Kawasaki and Seki [489], in the LAMP-Net architecture, use a vector map to define lane geometry and introduces a novel lane feature that encodes the geometric relationship between a vehicle’s state and surrounding lanes, enabling lane-aware trajectory predictions. Similarly, Bahari et al. [490] extract lane geometry, centerlines, and boundaries from HD maps to generate scene-compliant initial trajectory predictions. Lastly, DESIRE [64] also integrates map-related knowledge through the Scene Context Fusion (SCF) layer, which utilizes static semantic information from maps, such as road layouts, drivable areas, and road boundaries, to provide contextual guidance for trajectory generation.
5 Accounting for Uncertainty
@@
External uncertainty arises from the environment and includes factors beyond the model’s control. These encompass unobservable agent intentions, occlusions that limit visibility, environmental variability such as adverse lighting or weather conditions, and unexpected objects or rare events, including road accidents. For example, violations of traffic rules, such as crossing solid lanes or running red lights, can severely impact prediction performance by expanding the set of plausible future trajectories and degrading model reliability. Another challenge arises from rare events, like animals crossing the road are often underrepresented in training data, leading to potential failures in scenarios that deviate from typical urban driving conditions.
Internal uncertainty stems from limitations within the prediction system itself. It reflects challenges such as limited generalization to unseen scenarios, biases in the training data, and restricted model capacity to fully capture the complexity of real-world driving behavior. This form of uncertainty exists even when the inputs are ideal, due to incomplete knowledge or overly rigid modeling assumptions. In modular pipelines, one common source of internal uncertainty is propagated error — when imperfections in earlier modules affect downstream predictions. For instance, in the detect–track–predict setup, inaccurate detections or tracking failures can distort the input to the prediction module. However, most models are trained on clean, ground-truth trajectories and assume ideal upstream performance. In practice, this assumption breaks down, and prediction models must be designed to handle noisy inputs or account for upstream errors explicitly. More generally, any system that relies on outputs from other preceding modules is vulnerable to accumulated uncertainty along the pipeline.
@@
By nature, uncertainty can be categorized into aleatoric and epistemic types. Aleatoric uncertainty arises from intrinsic randomness in the data. In the context of autonomous driving, it is often linked to sensor characteristics, such as measurement noise, limited sensing range, calibration errors, or low resolution. These factors introduce ambiguities that pose inherent challenges to prediction models. Aleatoric uncertainty reflects the inherent difficulty in fully capturing the environment. For instance, predicting whether a pedestrian will cross at a crosswalk or continue walking along the sidewalk involves uncertainty due to the multimodal nature of human behavior. Aleatoric uncertainty can be further categorized as heteroscedastic (varying across samples, e.g., lighting conditions affecting image quality) or homoscedastic (uniform across samples, e.g., consistently miscalibrated sensors). Epistemic uncertainty, on the other hand, arises from limited knowledge of the environment, often due to insufficient or non-diverse training data. It can be reduced through improved data collection, better modeling techniques, and increased exposure to diverse scenarios. For example, cross-dataset domain adaptation is critical for addressing epistemic uncertainty. Many motion prediction models assume that training, testing, and deployment data share the same distribution, but this assumption rarely holds in real-world applications. The resulting domain gap can degrade performance, especially when encountering rare or previously unseen scenarios. Addressing epistemic uncertainty is therefore essential for improving model generalization and ensuring reliable operation in diverse, real-world settings.
Accounting for different sources of uncertainty and designing uncertainty-aware prediction models is an active area of research. Quantifying and mitigating uncertainty is essential for reducing associated risks and ensuring the safe operation of autonomous vehicles. Figure 11 summarizes the categorization of uncertainty based on its sources and nature. This section further discusses techniques for both mitigating and quantifying uncertainty. For a more detailed examination of uncertainty sources and mitigation approaches in autonomous driving, we refer the reader to the review by Yang et al. [492].
@@ 5.1 Mitigating Uncertainty
The effectiveness of data-driven prediction models heavily depends on the quality and diversity of their training datasets. Out-of-distribution (OOD) scenarios, where traffic conditions differ substantially from the training data, lead to distributional shifts that challenge prediction models during real-world operations. To mitigate this uncertainty, it is crucial to ensure that the training data provides substantial coverage of diverse scenarios and conditions. However, defining what constitutes “substantial coverage” remains an open research question. Laurent et al. [500] proposed a criterion for comprehensive testing, requiring scenarios to cover key parameters of autonomous driving systems that affect decision-making, while employing statistical methods to manage uncertainties during simulation runs.
To address uncertainties from incomplete inputs, Liao et al. [501] proposed a physics-informed model employing a Wavelet Reconstruction network to recover missing historical data and a Kinematic Bicycle model to enforce physically plausible predictions. A Risk-Aware module further integrates real-time risk factors using probabilistic attention, enabling adaptation to underrepresented scenarios. Bhattacharyya et al. [502] mitigated aleatoric and epistemic uncertainty by employing synthetic likelihoods, which encourage diverse model predictions and better capture multi-modal data. Synthetic likelihoods relax the requirement to explain every data point, enabling the model to generate diverse hypotheses for uncertain and multi-modal scene predictions. Next, Zhang et al. [503] introduced the Out-of-Sight Prediction framework (OOSTraj), which mitigates uncertainty in predictions for out-of-sight agents by utilizing a mobile denoising encoder and a camera parameters estimator to address sensor noise and map sensor-based to visual trajectories. The framework employs a Transformer-based decoder to predict trajectories even in the absence of direct visual observations.
Pourkeshavarz et al. [504] proposed a causal disentanglement approach to mitigate uncertainty caused by distribution shifts. Using a Transformer-GNN with disentangled attention blocks, the model separates causal and spurious factors in driving scenes, simulates potential shifts, and incorporates an invariance objective to focus on causal features while suppressing spurious ones for enhanced robustness. Finally, Luo et al. [40] addressed cascaded uncertainty by designing an end-to-end framework, “Fast and Furious” (FaF), which integrates 3D detection, tracking, and motion forecasting into a single convolutional network. By jointly optimizing these tasks, FaF implicitly mitigates cascading errors across the detection and prediction pipeline.
@@ 5.2 Quantifying Uncertainty
To minimize driving risks, it is essential to quantify motion prediction uncertainty and integrate this information into the decision-making module for safer and more robust planning. A common approach is to generate multi-modal predictions for each agent [248, 483, 119, 61, 505], rather than relying on a single deterministic trajectory. Probabilistic predictions address the uncertainty caused by unknown intentions of traffic participants, where a single past trajectory can unfold into multiple plausible futures. Some approaches estimate distribution parameters to provide a probabilistic representation of potential outcomes, such as modelling future trajectories using Gaussian distributions [506, 113, 507, 508] or mixture models [61, 122]. Table 15 summarizes deep learning-based models for quantifying uncertainty.
@@ 6 Planning-conditioned Trajectory Prediction
Planning-conditioned trajectory prediction methods rely on the assumption that agents act rationally and that their future trajectories are guided by a motion planner. Recent SOTA models [511, 512, 513, 514] effectively adopt this paradigm by introducing planning-prediction-coupled frameworks, where predictions are conditioned on the observed agents’ motion plans or intentions. These methods go beyond solely relying on past motion, contextual, and interaction features; by modelling the notion that agents follow potentially planned trajectories or have specific destinations to achieve.
To review planning-conditioned approaches, we follow two categorizations. First, these approaches are divided into model-based planning, learning-based planning, and combined methods, which are reviewed in Section 6.1. The second categorization focuses on anchor-based prediction methods, where trajectories are generated based on predefined anchors or guidance points derived from the environment or learned from data. These anchors typically represent potential future goals, paths, or key waypoints, that influence trajectory prediction. Anchor-based methods can be further divided into goal anchor-based prediction, where tentative goals for the observed agent are estimated online, followed by the generation of paths toward those goals, and trajectory-anchor-based prediction, where trajectories are directly constructed based on environmental or physical constraints. This categorization is further explored in Section 6.2.
@@ 6.1 Planning Methods for Trajectory Prediction
Motion planning literature has extensively developed methods for generating safe and efficient trajectories for ego vehicles. This has inspired the application of planning algorithms to predict the trajectories of other agents, under the assumption that their future paths align with specific objectives. Accordingly, planning-conditioned trajectory prediction methods adopt a similar categorization to motion planning: model-based, learning-based, and combined approaches [515]. Model-based approaches rely on explicit rules, physical constraints, and optimization techniques, often using kinematic or dynamic models to ensure physical plausibility. In contrast, learning-based approaches infer plans or intentions directly from observed data. Combined methods aim to integrate the strengths of both paradigms. Reviewed works using these approaches for trajectory prediction are summarized in Table 16.
@@ 6.1.1 Model-based planning approaches
Model-based planning for trajectory prediction forecasts agents’ future paths by leveraging models that simulate dynamics, kinematic constraints, decision-making processes, and interactions with the environment. These approaches integrate physical motion principles, behavioral models, control objectives, and road structure information to generate interpretable and feasible trajectories that reflect agents’ capabilities [520, 521]. Core components include: dynamics models representing how agents’ state evolves over time based on kinematic equations or physical motion laws [522]; cost functions, which define optimization objectives such as minimizing energy use, travel time, or collision risks [516]; and planners used to solve the resulting optimization problems under physical and contextual constraints.
@@ 6.1.2 Learning-based planning approaches
Learning-based planning for trajectory prediction employs data-driven models to generate trajectories for observed dynamic agents [511]. These methods often use reinforcement learning to infer planning strategies directly from data. By training on large datasets, they capture both high-level decision-making (e.g., navigating junctions or changing lanes) and low-level control behaviors (e.g., reacting to obstacles). Unlike traditional model-based methods that depend on explicit motion models, learning-based planning adapts to diverse, unstructured environments by learning implicit behavioral patterns. Moreover, by modeling multi-agent interactions, these approaches produce robust predictions in dynamic and complex traffic scenarios [61, 408, 64].
Previtali et al. [104] employ IRL to model agent behavior by inferring cost functions from observed trajectories. IRL is used to predict pedestrian trajectories by modelling their interactions with the environment, such as avoiding collisions or heading toward crosswalks. The approach involves learning a cost function that represents human decision-making processes and using a planning algorithm to generate plausible future paths based on these inferred behaviors. Further, the study by Deo and Trivedi [248] assumes that each surrounding agent can perform one of several maneuvers, such as lane changes or braking, which are learned using LSTMs, and predicts the corresponding trajectories by assigning probabilities to these maneuver classes.
Next, Ziebart et al. [523] proposed a model that assumes pedestrians behave as rational planners, efficiently moving toward their destinations. The prediction model learns a cost function that explains observed pedestrian trajectories and uses it to predict future positions. The ego vehicle’s planning module integrates this predictive model to generate a trajectory that avoids pedestrian obstacles while pursuing its own objectives. The intention-interaction graph method proposed by Li et al. [524] jointly models pedestrians’ intentions (goals) and their social interactions. The graph represents each pedestrian and their intended destination as nodes, enabling the model to predict trajectories by inferring goals based on observed behaviors and interactions, without requiring explicit knowledge of the final destinations.

### v3

@@ 4.2 Injecting Driving Knowledge
Table 14 provides a brief summary of learning-based models that incorporate driving knowledge, including the type of knowledge and the way it is integrated.
@@
For agent-related knowledge, a set of works focus on integrating kinematic parameters. Xie et al. [231] proposed the Interactive Multiple Model (IMM) method leveraging vehicle kinematic parameters such as acceleration and velocity. Cui et al. [495] embedded a kinematic model of two-axle vehicles into the prediction pipeline, constraining outputs like turning radius, acceleration, and heading to align with vehicle dynamics. Additionally, works [497, 488] incorporated vehicle dynamics and kinematics constraints through Model Predictive Control (MPC), ensuring that predicted trajectories comply with speed limits and acceleration.
@@
In terms of traffic-related knowledge, DESIRE [66] incorporates this knowledge through a predefined cost function within its Inverse Optimal Control (IOC)-based ranking module. The cost function evaluates trajectory predictions based on criteria such as collision avoidance, road compliance, and interaction dynamics, assigning lower costs to socially and contextually plausible trajectories to ensure alignment with expected traffic norms. Greer et al. [493] proposed the YawLoss framework, which integrates traffic-related knowledge by enforcing vehicle motion constraints aligned with lane directionality. This is achieved through a designed auxiliary loss function (YawLoss), which penalizes trajectories deviating from the expected lane heading ensuring lane-following behavior. Cho et al. [322] integrates traffic-related knowledge by representing predefined traffic rules as Signal Temporal Logic (STL) formulas. The model jointly predicts future trajectories and the degree of satisfaction (robustness slackness) for these rules, filtering out invalid trajectory predictions.
@@
Map-related knowledge represents the largest class of models integrating driving knowledge Bahari et al. [497], Lee et al. [66]. Wang et al. [498] proposed a lane-based prediction model that represents lane segments as vectorized elements, encoding geometric and semantic information and enabling interaction modelling between agents and lane segments. Yoon et al. [494] proposed a road-aware prediction model that incorporates curvilinear coordinates, lane boundaries, and feasible maneuvers from HD maps, using this information as prior knowledge to constrain the predictions. Next, Kawasaki and Seki [496] introduced the LAMP-Net architecture for lane-aware trajectory prediction, by encoding the geometric relationship between a vehicle’s state and surrounding lanes. Lastly, DESIRE [66] integrates map-related knowledge through the Scene Context Fusion (SCF) layer, which utilizes static semantic information from maps, such as road layouts, drivable areas, and road boundaries, to provide contextual guidance for trajectory generation.
5 Uncertainty Estimation and Mitigation
@@
Accounting for different sources of uncertainty and designing uncertainty-aware prediction models is an active research area across both academia and industry. For robust autonomous prediction, uncertainty must be explicitly modeled and systematically addressed at each stage of the autonomy stack. This section provides a high-level overview to build intuition on what constitutes uncertainty-aware modeling and why such awareness is critical. It covers sources of uncertainty (please refer to Figure 11), techniques for both quantifying and mitigating uncertainty. However, we strongly recommend readers interested in a deeper treatment of uncertainty across the full autonomy pipeline to refer to the reviews by Yang et al. [499], Wang et al. [152].
5.1 Sources of Uncertainty
@@
External uncertainty arises from the environment and includes factors beyond the model’s control. These encompass unobservable agent intentions, occlusions that limit visibility, environmental variability such as adverse lighting or weather conditions, and unexpected objects or rare events, including road accidents. For example, violations of traffic rules, such as crossing solid lanes or running red lights, can severely impact prediction performance by expanding the set of plausible future trajectories and degrading model reliability. Another challenge arises from rare events, like animals crossing the road are often underrepresented in training data.
Internal uncertainty arises from limitations within the prediction system itself. It reflects factors such as poor generalization to unseen scenarios, biases in the training data, and insufficient model capacity to capture the complexity of real-world driving behavior. A common contributor to internal uncertainty is propagated error, which originate from upstream modules that affect downstream predictions. For example, perception systems may produce imprecise outputs, such as shifted or shrunken bounding boxes, or entirely miss detections. When prediction is tightly coupled with perception, as in joint learning setups, such errors directly degrade forecasting performance. When connected through a tracker, perception errors may be partially compensated by estimations, but this can introduce new issues such as ID switches or misassociations, resulting into perturbed past trajectories. Consequently, any system that relies on outputs from earlier modules is inherently susceptible to accumulating uncertainty throughout the pipeline.
@@
By nature, uncertainty can be categorized into aleatoric and epistemic. Aleatoric uncertainty arises from intrinsic randomness in the data. In the context of autonomous driving, it is often linked to sensor characteristics, such as measurement noise, limited sensing range, calibration errors, or low resolution. Aleatoric uncertainty reflects the inherent difficulty in fully capturing the environment. For instance, predicting whether a pedestrian will cross at a crosswalk or continue walking along the sidewalk involves uncertainty due to the multimodal nature of human behavior. Aleatoric uncertainty can be further categorized as heteroscedastic (varying across samples, e.g., lighting conditions affecting image quality) or homoscedastic (uniform across samples, e.g., consistently miscalibrated sensors). Epistemic uncertainty, on the other hand, arises from limited knowledge of the environment, often due to insufficient or non-diverse training data. It can be reduced through improved data collection, better modeling techniques, and increased exposure to diverse scenarios, e.g., cross-dataset domain adaptation. Many motion prediction models assume that training, testing, and deployment data share the same distribution, but this assumption rarely holds in real-world applications. The resulting domain gap can degrade performance, especially when encountering rare or previously unseen scenarios. Addressing epistemic uncertainty is therefore essential for improving model generalization and ensuring reliable operation in diverse, real-world settings.
@@ 5.2 Mitigating Uncertainty
The reliability of data-driven trajectory prediction models heavily depends on the representativeness of the training data. When deployed in out-of-distribution (OOD) scenarios, where traffic conditions deviate significantly from those seen during training, model performance often degrades due to distributional shifts. One challenge lies in determining whether the training data sufficiently captures the variability of real-world environments. To this end, Laurent et al. [507] propose a structured testing framework, advocating that scenario coverage should span key factors influencing autonomous decision-making, with statistical tools used to handle simulation uncertainties. Beyond dataset diversity, models should also be equipped with mechanisms to detect OOD inputs at runtime. OOD detection methods include label-based techniques that monitor prediction error, Bayesian approaches that estimate epistemic uncertainty by modeling distributions over outputs or parameters [508, 509], and generative techniques that learn the input data distribution and flag deviations [510].
To address uncertainty sourced from incomplete inputs, Liao et al. [511] proposed a physics-informed model employing a Wavelet Reconstruction network to recover missing historical data and a Kinematic Bicycle model to enforce physically plausible predictions. A Risk-Aware module further integrates real-time risk factors using probabilistic attention, enabling adaptation to underrepresented scenarios. Bhattacharyya et al. [512] mitigated aleatoric and epistemic uncertainty by employing synthetic likelihoods, which relax the requirement to explain every data point, enabling the model to generate diverse hypotheses for uncertain and multi-modal cases. Next, Zhang et al. [513] introduced the Out-of-Sight Prediction framework (OOSTraj), which mitigates uncertainty in predictions for out-of-sight agents by utilizing a mobile denoising encoder and a camera parameters estimator to address sensor noise and map sensor-based to visual trajectories. The framework employs a Transformer-based decoder to predict trajectories even in the absence of direct visual observations.
Pourkeshavarz et al. [514] proposed a causal disentanglement approach to mitigate uncertainty caused by distribution shifts. Using a Transformer-GNN with disentangled attention blocks, the model separates causal and spurious factors in driving scenes, simulates potential shifts, and incorporates an invariance objective to focus on causal features while suppressing spurious ones for enhanced robustness. Next, Luo et al. [42] addressed cascaded uncertainty by designing an end-to-end framework, “Fast and Furious” (FaF), which integrates 3D detection, tracking, and motion forecasting into a single model. By jointly optimizing these tasks, FaF implicitly mitigates cascading errors across the detection and prediction pipeline.
@@ 5.3 Quantifying Uncertainty
Quantifying prediction uncertainty and incorporating it into downstream decision-making can reduce driving risks by enabling more cautious and informed planning. Instead of relying on a single deterministic forecast, many models generate multi-modal predictions for each agent [261, 490, 122, 63, 515]. This helps capture uncertainty stemming from the inherent ambiguity in agents’ future intentions, since a single past trajectory can lead to multiple plausible futures. To represent these possibilities, some methods estimate probability distributions over future trajectories. Common approaches include modeling outcomes using parametric forms such as Gaussian distributions [516, 116, 517, 518] or Gaussian Mixture Models (GMMs) [63, 128], which allow capturing both the mean prediction and its associated uncertainty. Table 15 summarizes deep learning-based models for quantifying uncertainty.
@@ 6 Planning-conditioned Trajectory Prediction
Planning-conditioned trajectory prediction methods rely on the assumption that agents act rationally and their future trajectories are guided by a motion planner. Models [521, 522, 523, 524] effectively adopt this paradigm by introducing planning-prediction-coupled frameworks, where predictions are conditioned on observed motion plans or intentions. These methods go beyond solely relying on past trajectory, contextual, and interaction features; by modelling the notion that agents follow potentially planned trajectories or have specific destinations to reach.
To review planning-conditioned approaches, we follow two categorizations. In the first category, the approaches are divided into model-based planning, learning-based planning, and combined methods, which are reviewed in Section 6.1. In the second category, it focuses on anchor-based prediction methods, where trajectories are generated based on predefined anchors or guidance points derived from the environment or learned from data. These anchors typically represent potential future goals, paths, or key waypoints. Anchor-based methods can be further divided into goal anchor-based prediction, where tentative goals for the agents are estimated, and trajectory-anchor-based prediction, where possible trajectories are directly constructed based on environmental and physical constraints. This categorization is further explored in Section 6.2.
@@ 6.1 Planning Methods for Trajectory Prediction
Motion planning literature has extensively developed methods for generating safe and efficient trajectories for ego vehicles. This has inspired the application of planning algorithms to predict the trajectories of other agents, under the assumption that their future paths align with specific objectives. Accordingly, planning-conditioned trajectory prediction methods adopt a similar categorization to motion planning: model-based, learning-based, and combined approaches [525]. Model-based approaches rely on explicit rules, physical constraints, and optimization techniques. In contrast, learning-based approaches infer plans or intentions directly from observed data. Combined methods aim to integrate the strengths of both paradigms. Reviewed works using these approaches for trajectory prediction are summarized in Table 16.
@@ 6.1.1 Model-based planning approaches
Model-based planning approaches integrate physical motion principles, behavioral models, control objectives, and road structure information to generate interpretable and feasible future trajectories [530, 531]. Core components include: dynamics models representing how agents’ state evolves over time based on kinematic equations or physical motion laws [532]; cost functions, which define optimization objectives such as minimizing energy use, travel time, or collision risks [526]; and planners used to solve the resulting optimization problems under physical and contextual constraints.
@@ 6.1.2 Learning-based planning approaches
Learning-based planning for trajectory prediction employs data-driven models to generate trajectories for target agents [521]. These methods often employ reinforcement learning to infer planning strategies directly from data. By training on large datasets, they capture both high-level decision-making (e.g., navigating junctions or changing lanes) and low-level control behaviors (e.g., reacting to obstacles). By learning implicit behavioral patterns and modeling multi-agent interactions, these approaches generate predictions in dynamic and complex traffic scenarios [63, 417, 66].
Previtali et al. [107] use inverse reinforcement learning (IRL) to predict pedestrian trajectories by inferring cost functions from observed behavior. The learned cost captures decision-making patterns such as avoiding collisions or moving toward crosswalks, and is used in a planner to generate plausible future paths. Further, the study by Deo and Trivedi [261] uses LSTM to predict maneuver classes (e.g., lane changes, braking) for target agents, and generates corresponding trajectories by assigning probabilities to each maneuver. Next, model [533] learns a cost function from observed pedestrian trajectories to predict future positions. The ego vehicle plans its trajectory by jointly considering predicted pedestrian motion and its own objectives. The intention-interaction graph method proposed by Li et al. [534] jointly models pedestrians’ intentions (goals) and their social interactions. The graph represents each pedestrian and their intended destination as nodes, enabling the model to predict trajectories by inferring goals based on observed behaviors and interactions.
In summary, learning-based planning for trajectory prediction leverages large-scale, real-world data to uncover both strategic and reactive driving patterns. By using reinforcement and inverse reinforcement learning, these approaches infer cost functions and planning policies directly from observed behavior, enabling the generation of realistic future paths for both vehicles and pedestrians